Circular No. 8575

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

NOVA SCORPII 2005 M. L. Sitko and R. Kimes, University of Cincinnati; D. K. Lynch, R. W.

Russell, and D. L. Kim, The Aerospace Corporation; and D. Griep, Infrared Telescope Facility (IRTF), report 3–14-μm spectroscopy of this object on July 27.375 UT using the instrument BASS on the IRTF. The spectrum shows a smooth, featureless continuum consistent with a Rayleigh-Jeans tail of a Planck function. The magnitude was ≈ 4.8 in the bands L, M, and N.

Visual magnitude estimates: July 26.735 UT, 8.9 (T. Cooper, Bredell, Kempton Park, S. Africa); 26.881, 8.9 (J. G. de S. Aguiar, Campinas, Brazil); 27.228, 9.0 (M. Linnolt, Woodside, CA).

SUPERNOVA 2005dc IN NGC 7107 P. Luckas, Perth, W. Australia; O. Trondal, Oslo, Norway; and M.

Schwartz, Patagonia, AZ, report the discovery of an apparent supernova (mag 15.7) on an unfiltered CCD frames taken with a 0.35-m Tenagra telescope at Perth on July 26.70 and 27.62 UT in the course of the Tenagra Observatory Supernova Search. SN 2005dc is located at α = 21^{h}42^{m}22^{s}.66, δ = −44^{o}47′21′′.2 (equinox 2000.0), which is 41′′ west and 9′′.6 north of NGC 7107 (the galaxy center having position end figures 26^{s}.50, 30′′.8). Nothing was visible at the position of 2005dc on an image taken by Luckas on July 5.61 (limiting magnitude ∼ 18.5).

SUPERNOVA 2005da IN UGC 11301 M. Modjaz, P. Challis, and R. Kirshner, Harvard-Smithsonian Center

for Astrophysics; and T. Matheson, National Optical Astronomy Observatory, report that a spectrogram (range 350–740 nm) of SN 2005da (cf. IAUC 8570), obtained by P. Berlind on July 26.26 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST), reveals it to be a supernova, possibly a peculiar type-Ic supernova. No obvious lines of hydrogen or helium are detected. The spectrum consists of wide absorption features, similarly to peculiar type-Ic supernovae 1998bw and 2002ap (Foley et al. 2003, PASP 115, 1220). In contrast to the spectra of supernovae 1998bw and 2002ap, this new object’s spectrum also exhibits a wide absorption trough at 585 nm, having removed the NED recession velocity of 4495 km/s for the host galaxy (from Falco et al. 1999, PASP 111, 438), which could be identified with Na I. There are peaks near 530 and 640 nm and absorptions near 470, 630, and 680 nm.

2005 July 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green